Fermi National Laboratory
 

Accelerator Update for January 7, 2009 - January 9, 2009

 

Wednesday Jan 7

The day shift began with Operations monitoring the Tevatron (TeV) store 6707 (114.61E30) and the Recycler stash (200.6E10), with the Antiproton Source stacking (34.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam.

At 9:20 AM, two Booster RF stations (BRF15 & 16) tripped off. They reset.

At 9:25 AM, BRF7 tripped off. It reset.

At 9:28 AM, BRF15 tripped off. When operators reset the station, BRF11 tripped off. It reset. These RF stations were tripping off due to tuning by experts.

The Neutron Therapy Facility reported at 10:08 AM that they had completed treating their last patient for the day.

At 10:27 AM, the P1 transfer line beam permit tripped off due to low current on a power supply (V714). Everything reset.

At 10:29 AM, BRF11 & 16 tripped off. They reset.

Operations reduced the beam intensity to MiniBooNE at 10:37 AM in preparation for an access later in the day.

At 10:55 AM, BRF7 tripped off for the first of two times in a minute. It reset.

At 10:56 AM, BRF11 tripped off for the first of three times in 18 minutes. It reset.

At 11:01 AM, the Booster East Anode Power Supply tripped off due to a trip by BRF5. Everything reset.

At 11:14 AM, the 400 MeV area beam permit tripped off for the first of two times in about an hour due to tuning by a Linac expert. It reset.

At 11:20 AM, BRF4 tripped off. It reset.

At 12:08 PM, BRF9 & 11 tripped off. They reset.

At 12:15 PM, MIRF14 tripped off. It reset.

At 12:36 PM, BRF3 & 6 tripped off for the first of two times in six minutes. They reset.

Operations put Pbar into standby at 12:37 PM due to the ongoing low-level RF work on LRF5, which was disrupting beam.

Operations terminated store 6707 at 1:03 PM.

Operations prepared CDF and D0 for access at 1:22 PM.

Cryo system technicians began their work on the TeV sectors A3 and D1 wet engines at 1:24 PM.

Booster experts reported at 1:39 PM that they found their RF problem; an amplifier in the low level RF room had failed. It was replaced.

D0 came out of access at 2:12 PM.

Pbar resumed stacking at 2:39 PM.

At 2:51 PM, a Linac Klystron RF station (KRF5) tripped off and BRF7 tripped off. They both reset.

At 3 PM, the 400 MeV area tripped off. It reset.

Operations disabled the MiniBooNE beam at 3:04 PM for a MI power supply (MBEX) study and work.

At 3:19 PM, KRF3 tripped off. It reset.

Cryo system techs reported at 3:21 PM that they had completed their wet engine work.

Operations prepared MiniBooNE for access at 3:39 PM.

CDF came out of access at 3:49 PM.

Operations reported at 3:57 PM that the Recycler had stashed antiprotons five times during the day shift.

The MBEX power supply work ended at 4:29 PM.

Operations put the TeV through a dry squeeze at 4:32 PM.

At 4:51 PM, BRF9 & 11 tripped off. They reset.

At 4:55 PM, BRF4 tripped off. It reset.

At 4:58 PM, Operations brought the TeV down to 150 GeV. Operators discovered a power supply (VB25) that increased its current output until it tripped off. To fix the problem, they replaced the Digital Function Generator chassis.

At 5:02 PM, BRF3 & 4 tripped off. They reset.

At 5:15 PM, BRF4 tripped off. It reset.

Operations put the TeV through a wet squeeze at 5:34 PM.

Operations put the TeV into shot setup around 6:14 PM.

A NuMI expert reported at 7:04 PM that they had removed the OTR carbon foil from the beam path.

Operations prepared Booster for access at 7:09 PM.

Operations put Pbar into standby at 7:12 PM.

Operations prepared MI for access at 7:33 PM.

LCW experts at CUB reported at 7:49 PM that they were bringing column #1 online.

Operations established store 6710 at 7:52 PM with an initial luminosity of 330.0E30.

FESS experts reported at 7:55 PM that they had completed their air handling work at AP0.

Booster came out of access at 8:08 PM.

Operations established beam through Linac and Booster at 8:39 PM.

MI came out of access at 10:24 PM.

Operations established beam to MiniBooNE at 10:40 PM.

Operations established beam to NuMI and Pbar at 11:06 PM.

Operations reported at midnight that the Recycler had stashed four times during the evening shift.

Thursday Jan 8

The midnight shift began with Operations monitoring store 6710 (162.94E30) and the stash (31.9E10), with Pbar stacking (20.8E10), and with NuMI and MiniBooNE taking beam.

At 12:13 AM, MIRF14 tripped off and would not reset. Operators bypassed the station.

At 1:08 AM, the P1 transfer line beam permit tripped off when a power supply’s (V714) output current drifted low. Everything reset.

At 1:43 AM, Operations reported that a MiniBooNE expert arrived to help troubleshoot autotune problems.

The MiniBooNE expert reported at 2:59 AM that he found a BPM with reversed wiring.

At 4:17 AM, V714 caused the P1 transfer line to trip. Everything reset.

At 6:09 AM, a Linac quadrupole power supply’s (QPS232) output current began to climb and then went unstable. Operators made adjustments to compensate until the supply went stable, which was about an hour later.

Operations reported at 8:01 AM that the Recycler had stashed eight times during the midnight shift.

At 8:09 AM, LRF4 tripped off. It reset.

At 8:13 AM, Operations reported that an Accumulator RF station (ARF1) had tripped off. An expert investigated.

Operations put Pbar into standby at 8:26 AM.

Pbar resumed stacking at 8:37 AM.

At 9:53 AM, MIRF10 tripped off for the first of two times in three minutes. It reset.

At 10:59 AM, the P1 transfer line tripped off due to V714. Everything reset.

At 12:59 PM, BRF6 & 7 tripped off. They reset.

Operations terminated store 6710 at 1:03 PM.

Cryo system techs reported at 1:14 PM that they were beginning to adjust connecting rod bearings for wet engines at TeV sectors A4, B2, C4, and E4.

Operations prepared D0 and CDF for access at 1:27 PM.

At 1:45 PM, BRF7, 9, and 16 tripped off. They all reset.

At 1:46 PM, BRF7 & 9 tripped off. They reset.

At 1:51 PM, BRF8 tripped off for the first of three times in 24 minutes.

At 1:59 PM, BRF6 tripped off for the first of two times in 13 minutes.

MI experts began a coalescing study at 2:03 PM.

At 2:13 PM, BRF3, 4, 7, and 9 tripped off. They all reset.

At 2:14 PM, BRF7 & 11 tripped off for the first of three times in one minute. They reset. Booster experts began investigating all these spark-induced trips.

CDF came out of access around 2:27 PM.

Cryo system techs reported that they had completed their wet engine work at 2:49 PM.

At 3:12 PM, three Booster kickers (KPS1A, B, and C) tripped off. They all reset.

Operations reported at 4 PM that the Recycler had stashed seven times during the day shift.

D0 came out of access around 4:25 PM.

Operations put the TeV through a dry squeeze at 4:51 PM.

Operations put the TeV through a wet squeeze at 5:26 PM.

Operations put the TeV into shot setup at 5:54 PM.

Operations established store 6713 at 7:37 PM with an initial luminosity of 329.9E30.

Operations reported at midnight that the Recycler had stashed six times during the evening shift.

Friday Jan 9

The midnight shift began with Operations monitoring store 6713 (161.47E30) and the stash (135.5E10), with Pbar stacking (36.1E10), and with NuMI and MiniBooNE taking beam.

At 3:06 AM, LRF5 tripped off for the first of two times in five minutes. It reset.

Around 5:14 AM, Operations reported a problem with stacking. TWTs were tripping off. Operators investigated. Stacking continued, but at a reduced rate.

A Pbar expert began to help investigate the stacking problem at 6:07 AM.

At 6:44 AM, a Linac quadrupole power supply (QPS111) tripped off. It reset.

Another Pbar expert reported 7:50 AM that the stacking problem was due to a PLC in the tunnel. He reported that an access would be needed to make repairs.

Operations reported at 7:57 AM that the Recycler had stashed eight times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Pbar experts to access their tunnel this afternoon to fix their stacking problem.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/09/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory